# revtex4-2Repair the float

### v1

@@ Abstract
Hyperbolic models can reproduce the heavy-tailed degree distribution, high clustering, and hierarchical structure of empirical networks. Current algorithms for finding the hyperbolic coordinates of networks, however, do not quantify uncertainty in the inferred coordinates. We present BIGUE, a Markov chain Monte Carlo (MCMC) algorithm that samples the posterior distribution of a Bayesian hyperbolic random graph model. We show that combining random walk and random cluster transformations significantly improves mixing compared to the commonly used and state-of-the-art dynamic Hamiltonian Monte Carlo algorithm. Using this algorithm, we also provide evidence that the posterior distribution cannot be approximated by a multivariate normal distribution, thereby justifying the use of MCMC to quantify the uncertainty of the inferred parameters.
@@ I Introduction
Embedding enables us to comprehend abstract objects and conduct rigorous quantitative analyses of their similarities, differences, and structural characteristics. In the complex network context, embeddings usually consist of node coordinates in a latent space. A rich literature has shown that embedding in hyperbolic spaces [1], in particular, can visually emphasize the important parts of large networks [2], but also, when used as the basis for a random graph model, naturally reproduce many common network properties such as the community structure [3, 4, 5, 6, 7], a power-law degree distribution [8, 9], a non-vanishing clustering coefficient [8, 10, 9, 11, 12], a fractal structure [13, 14, 15] and the sparsity of connections [12]. Furthermore, since hyperbolic geometry closely approximate the shortest paths of graphs via geodesics [8, 16], hyperbolic embeddings have been used to design efficient routing protocols [17, 18].
@@
Considerable effort has gone into representing a given graph using latent spaces [19, 20, 21, 22, 23]. Many studies focus on placing the vertices such that the graph distance between the vertices is closely approximated by the hyperbolic geodesics [24, 25, 26, 27, 28]. From this perspective, Ref. [29] proved that trees can be embedded in hyperbolic space up to an arbitrarily small error, enabling perfect greedy routing. This is due to the exponentially growing volume in the hyperbolic space, which can be related to the branching factor of trees [8]. The embedding task was also tackled with dimension reduction techniques from machine learning [23, 30, 31, 32] and with the maximum likelihood estimation of various hyperbolic random graph models [33, 17, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Other methods allow the embedding of directed graphs [43, 44, 45], weighted graphs [46, 47] and graphs to the $D$-dimensional hyperbolic space with [48] or without vertex features [49].
@@
Figure 1: Probabilistic hyperbolic embedding of a synthetic graph with BIGUE. Hyperbolic coordinates are obtained using a transformation between the $\mathbb{H}^{2}$ and the $\mathbb{S}^{1}$ models described in Appendix A. Black points and dark-colored points are the median coordinates of each vertex. Light-colored points are sampled positions for the three highlighted vertices. Lines are edges drawn using hyperbolic geodesics. We used $2,000$ samples corresponding to the simulations shown in Fig. 3d.
Existing methods have largely overlooked uncertainty: They return a single embedding without allowance for error or multimodality. Yet, this information is crucial as perturbations to the embedding affect link prediction [50] and can lead to different graph representations. As a first step towards addressing this problem, we present BIGUE (Bayesian Inference of a Graph’s Unknown Embedding), a Markov chain Monte Carlo (MCMC) algorithm that samples the posterior distribution of a Bayesian model constructed from the commonly used $\mathbb{H}^{2}$ hyperbolic random graph model (HRGM); see Fig. 1.
@@
The Bayesian hyperbolic embedding of networks has already been studied in Ref [51], but the results therein suggest that the proposed MCMC algorithm has poor mixing—a common problem for many standard and state-of-the-art MCMC methods when the posterior distribution is complex and multimodal [52, 53, 54, 55, 56, 57, 58, 59, 60]. We identify community structure [61] as a source of multimodality and introduce a set of cluster-based transformations to improve the exploration of the posterior distribution. Our approach draws on previous literature in which community structure was used to guide greedy embedding algorithms [41, 62]. We show that these transformations drastically improve the mixing of the MCMC algorithm, allowing us to find appropriate credible intervals for the parameters of the HRGM and any graph property. We also show that multimodality can be induced in the posterior by having a vertex with two positions and that it is identified by BIGUE.
II Bayesian hyperbolic embedding
II.1 Model definition
@@
Let $\mathcal{G}$ be a random graph and $G=(V,E)$ be the graph we aim to embed. We use the circular probabilistic model $\mathbb{S}^{1}$ to describe the graph $G$; it is nearly equivalent to the hyperbolic plane model $\mathbb{H}^{2}$ [8] (see Appendix A), but it will turn out to facilitate inference. In the $\mathbb{S}^{1}$ model, the probability that an edge $(u,v)$ exists between vertices $u$ and $v$ is a function of the angular coordinates $\theta=(\theta_{w})_{w\in V}$ of the vertices where each $\theta_{w}\in[-\pi,\pi)$, of the parameters $\kappa=(\kappa_{w})_{w\in V}$ where each $\kappa_{w}>0$ control the degree of vertices, and of an inverse temperature $\beta>1$ which controls the amount of clustering. This probability is defined as
$\mathbb{P}[(u,v)\in E|\theta,\kappa,\beta]=\quantity(1+\quantity(\frac{d(\theta_{u},\theta_{v})}{\mu\kappa_{u}\kappa_{v}})^{\beta})^{-1},$ (1)
@@
where $d(\theta_{u},\theta_{v})=R\Delta(\theta_{u},\theta_{v})$ is the arc length between vertices $u$ and $v$, $\Delta(\theta_{u},\theta_{v})=\pi-|\pi-|\theta_{u}-\theta_{v}||$ is the angular separation, and $R=|V|/2\pi$ is the radius of the circle on which vertices are embedded. We fix the constant $\mu=\beta\sin(\pi/\beta)/2\pi\mathbb{E}[\kappa]$ so that the parameters $\kappa$ match the degrees $\mathbb{E}[K(\kappa)]=\kappa$ in expectation in the limit of large graphs, where $K(\kappa)$ is the random degree of a vertex of parameter $\kappa$ when all vertices are independently and uniformly distributed on the circle [8].
@@
$\mathbb{P}[\mathcal{G}=G|\theta,\kappa,\beta]$ $=\prod_{(u,v)\in E}\mathbb{P}[(u,v)\in E|\theta,\kappa,\beta]$
×$\hskip 10.0pt\times\prod_{(u,v)\not\in E}(1-\mathbb{P}[(u,v)\in E|\theta,\kappa,\beta])$
@@
where $\binom{V}{2}$ is the set of all combinations of two vertices and $\beta_{uv}^{\pm}$ is $\beta$ if $(u,v)\in E$ and is $-\beta$ otherwise.
@@
The likelihood $\mathbb{P}[\mathcal{G}=G|\theta,\kappa,\beta]$ is given by Eq. (II.1). To complete the specification of Eq. (3), we use independent prior densities
@@
with $\beta_{0}=3$ and $\sigma=2$. Since known networks tend to have a heavy-tailed degree distribution, we use a half-Cauchy prior
@@
except for the two vertices with the highest degree, whose position is fixed (the reason is technical and has to do with avoiding degeneracy in the posterior; see below.) With these choices, the posterior distribution is proper and can express correlations between the parameters even if the priors are independent. The normalization constant $\mathbb{P}[\mathcal{G}=G]$ can be contained by integration in principle, though it is not needed here; see Sec. III below.
@@
Figure 2: Log-likelihood of the model when a single vertex is moved along the circle. Each dip is a divergence that occurs when a vertex is positioned at the same position as a non-adjacent vertex. These divergences form barriers in the objective function landscape (log-likelihood, posterior distribution), which is one of the reasons why hyperbolic embedding is a difficult problem.
II.2 Symmetries and automorphisms
Our Bayesian approach to hyperbolic embedding exacerbates two definitional issues that are usually ignored when the model is used with maximization algorithms.
@@ II.2 Symmetries and automorphisms
First, the model inherits the symmetries of the circle $\mathbb{S}^{1}$. Since the distance $d$ is invariant to rotations and reflections $\theta_{w}\mapsto-\theta_{w}$, the edge probabilities are preserved in the likelihood, and an infinite number of equivalent embeddings exist. Maximization algorithms break this symmetry at random, but we need to proceed with more caution when we move to a distribution over embedding. Without loss of generality, we fix the angular coordinate of the highest-degree vertex $u$ to $\theta_{u}=0$ to handle rotational symmetries. We remove the reflection symmetry by restricting the angular coordinate of the second highest-degree vertex $v$ to $\theta_{v}\in[0,\pi)$.
@@
Second, a graph is inherently labeled by its vertices, but any vertex relabelling that preserves the edges—an automorphism—leaves the likelihood invariant since the distances between connected and unconnected vertices are unchanged. For example, if a transformation $u\leftrightarrow v$ exists such that each $(u,w)\in E\implies(v,w)\in E$ and $(v,w)\in E\implies(u,v)\in E$, then we can also exchange $k_{u}\leftrightarrow k_{v}$ and $\theta_{u}\leftrightarrow\theta_{v}$ in the likelihood. We account for this symmetry by defining the likelihood up to an automorphism.
III Sampling
The posterior density, Eq. (3), does not correspond to the closed-form density of a known distribution, and computing analytical expectations thus appears difficult. Therefore, we turn to numerical estimates with sampling.
@@ III Sampling
Unfortunately, for most graphs $G$, the posterior distribution has a complicated geometry, making the application of well-known sampling algorithms challenging. This is the case even for small graphs, as Fig. 2 shows: changing the embedding coordinates of a single vertex reveals multiple local maxima of the likelihood (and thus posterior density) for a small graph of 30 vertices. The model predicts that two vertices with nearby embedding coordinates should be connected with high probability. In the extreme case where non-adjacent vertices have the same angular coordinate, the likelihood approaches 0, causing the sharp dips shown in the Figure. Further, the gradients are undefined due to their dependence on the distance function in Eq. (1). Since the posterior geometry is filled with these barriers for sparse and large graphs (the number of modes increases with the number of non-adjacent pairs of vertices), hyperbolic embedding can be challenging for standard algorithms.
@@
Figure 3: Statistics of Markov chain Monte Carlo algorithms when embedding a synthetic graph in $\mathbb{S}^{1}$. (a) Vertex and chain averaged autocorrelation for different lags. The autocorrelation of angular coordinates is adapted to consider their periodicity (see Appendix D). (b-c) Log-likelihood of a simulated Markov chain initialized with (panel b) and without (panel c) access to ground truth information. When ground truth is unknown, we initialize $\kappa_{u}=\deg(u)$ and draw the other parameters from their priors. (d) Log-likelihood of a Markov chain after thinning the chain shown in panel b. In each case, we compute $\hat{R}_{\text{max}}$, the maximum potential scale reduction factor for all parameters after the warm-up (grey) — values close to 1 are desirable. The graph has 30 vertices and is generated with the $\mathbb{S}^{1}$ likelihood, with angular coordinates drawn from their prior, $\beta=2.5$ and the $\kappa$ parameters drawn from a Pareto distribution of parameter $2.5$ and $\kappa_{\text{min}}=4$ and truncated to $10$. Four independent chains were simulated to compute $\hat{R}$ and the autocorrelation, but only one is displayed. The median effective sample sizes $S_{\text{eff}}$ across all parameters were, for BIGUE: $500$, $442$, and $1608$ (in panels b,c, and d respectively). For the random walk MCMC, we obtained medians of $S_{\text{eff}}=236$, and $411$, and for dynamic HMC, they were equal to $603$ and $3837$.
III.1 Baseline algorithms

### v2

@@ Abstract
Hyperbolic models are known to produce networks with properties observed empirically in most network datasets, including heavy-tailed degree distribution, high clustering, and hierarchical structures. As a result, several embeddings algorithms have been proposed to invert these models and assign hyperbolic coordinates to network data. Current algorithms for finding these coordinates, however, do not quantify uncertainty in the inferred coordinates. We present BIGUE, a Markov chain Monte Carlo (MCMC) algorithm that samples the posterior distribution of a Bayesian hyperbolic random graph model. We show that the samples are consistent with current algorithms while providing added credible intervals for the coordinates and all network properties. We also show that some networks admit two or more plausible embeddings, a feature that an optimization algorithm can easily overlook.
@@ I Introduction
Embedding enables us to comprehend abstract objects and conduct rigorous quantitative analyses of their similarities, differences, and structural characteristics. In the complex network context, embeddings usually consist of vertex coordinates in a latent space. A rich literature has shown that embedding in hyperbolic spaces [1], in particular, can visually emphasize the important parts of large networks [2], but also, when used as the basis for a random graph model, naturally reproduce many common network properties such as the community structure [3, 4, 5, 6, 7, 8, 9], a power-law degree distribution [10, 11], a non-vanishing clustering coefficient [10, 12, 11, 13, 14, 15], a fractal structure [16, 17, 18] and the sparsity of connections [14]. Furthermore, since hyperbolic geometry closely approximate the shortest paths of graphs via geodesics [10, 19], hyperbolic embeddings have been used to design efficient routing protocols [20, 21].
@@
Considerable effort has gone into representing a given graph using latent spaces [22, 23, 24, 25, 26]. Many studies focus on placing the vertices such that the graph distance between the vertices is closely approximated by the hyperbolic geodesics [27, 28, 29, 30, 31]. From this perspective, trees [32] can be embedded in hyperbolic space up to an arbitrarily small error, enabling perfect greedy routing. This is due to the exponentially growing volume in the hyperbolic space, which can be related to the branching factor of trees [10]. The embedding task was also tackled with dimension reduction techniques from machine learning [26, 33, 34, 35] and with the maximum likelihood estimation of various hyperbolic random graph models [36, 20, 37, 38, 39, 40, 41, 42, 43, 44, 45]. Other methods allow the embedding of directed graphs [46, 47, 48], weighted graphs [49, 50] and graphs to the $D$-dimensional hyperbolic space with [51] or without vertex features [52].
@@
Figure 1: Probabilistic hyperbolic embedding of a synthetic graph with BIGUE. Hyperbolic coordinates are obtained using a transformation between the $\mathbb{H}^{2}$ and the $\mathbb{S}^{1}$ models described in Supplementary Note 2. Black points and dark-colored points are the median coordinates of each vertex. Light-colored points are the $2000$ sampled positions for the three highlighted vertices. Lines are edges drawn using hyperbolic geodesics. The synthetic graph of 30 vertices is generated with the $\mathbb{S}^{1}$ likelihood, with angular coordinates drawn from their prior, $\beta=2.5$ and the $\kappa$ parameters drawn from a Pareto distribution of exponent $-2.5$ over the truncated interval $(4,10)$.
Existing methods have, however, largely overlooked uncertainty: They return a single embedding without allowance for error or acknowledging the existence of other solutions. Yet, this information is crucial as perturbations to the embedding affect edge prediction [53] and can lead to different graph representations.
As a first step towards addressing this problem, we propose a Bayesian embedding algorithm that quantifies uncertainty rigorously. Related methods have proven useful in network science, where they have been used to reconstruct networks from noisy measurements  [54, 55], predict the existence of edges [56], infer higher-order interactions from network data [57] and pairwise observation [58], and perform community detection [59, 60].
Our proposed algorithm, BIGUE (Bayesian Inference of a Graph’s Unknown Embedding), uses Markov chain Monte Carlo (MCMC) sampling to generate embeddings compatible with a given graph; see Fig. 1. We show how these samples can be used to calculate credible intervals, interpreted as error bars, for vertex positions, properties of the embedding, and model parameters. We also show that many qualitatively different embeddings are often compatible with an observed graph—a phenomenon known as multimodality.
@@
Bayesian approaches hyperbolic embeddings of networks have already been studied [61], but the results therein suggest that the proposed MCMC algorithm has poor mixing—a common problem many MCMC methods face when the posterior distribution is complex and multimodal [62, 63, 64, 65, 66, 67, 68, 69, 70]. We identify community structure [9] as a source of multimodality and introduce a set of cluster-based transformations to improve the exploration of the posterior distribution. Our approach draws on previous literature in which community structure was used to guide greedy embedding algorithms [44, 71]. We show that these transformations drastically improve the mixing of the MCMC algorithm, even if the community structure is weak. In doing so, we make a tradeoff between speed and accuracy—greedy methods find a single solution rapidly, while our MCMC methods explore a large space more slowly. In fact, our implementation of BIGUE does not scale well beyond a few hundred vertices. Thad said we show that the cluster transformations powering BIGUE improve mixing and thus make it more scalable than alternative MCMC algorithms. Furthermore, we provide evidence that the posterior distributions of hyperbolic graph models are not Gaussian in practice, thereby motivating the need for MCMC as opposed to faster but less accurate methods like variational inference.
II Methods
II.1 Bayesian hyperbolic embedding
The Bayesian framework aims to infer parameters—here the embedding—from some data—the graph. In this section, we define the likelihood and prior distributions that lead to the posterior of the model, and highlight unique challenges that do not arise when searching for a single embedding.
II.1.1 Model definition
@@
Let $G=(V,E)$ be the graph we aim to embed, with $V$ and $E$ being the sets of its vertices and of its edges, respectively. Let $\mathcal{G}$ be the random graph used to perform the embedding. We use the circular probabilistic model $\mathbb{S}^{1}$ to describe the graph $G$ [15]; it is nearly equivalent to the hyperbolic plane model $\mathbb{H}^{2}$ [10] (see Supplementary Note 2), but it will turn out to facilitate inference. In the $\mathbb{S}^{1}$ model, the probability that an edge $(u,v)$ exists between vertices $u$ and $v$ is a function of the angular coordinates $\theta=(\theta_{w})_{w\in V}$ of the vertices where each $\theta_{w}\in[-\pi,\pi)$, of the parameters $\kappa=(\kappa_{w})_{w\in V}$ where $\kappa_{w}>0$ controls the degree of vertex $w$, and of an inverse temperature $\beta>1$ which controls the amount of clustering. This probability is defined as
$\mathbb{P}[a_{uv}=1|\theta,\kappa,\beta]=\quantity(1+\quantity(\frac{d(\theta_{u},\theta_{v})}{\mu\kappa_{u}\kappa_{v}})^{\beta})^{-1},$ (1)
@@
where $a_{uv}=a_{vu}$, $u\neq v$, is any element of the adjacency matrix of a graph realized by $\mathcal{G}$, $d(\theta_{u},\theta_{v})=R\Delta(\theta_{u},\theta_{v})$ is the arc length between vertices $u$ and $v$, $\Delta(\theta_{u},\theta_{v})=\pi-|\pi-|\theta_{u}-\theta_{v}||$ is the angular separation, and $R=|V|/2\pi$ is the radius of the circle on which vertices are embedded. We fix the constant $\mu=\beta\sin(\pi/\beta)/(2\pi\mathbb{E}[\tilde{\kappa}])$ so that the parameters $\kappa$ match the degrees $\mathbb{E}[K|\kappa]=\kappa$ in expectation in the limit of large graphs, where $\tilde{\kappa}$ represents the random variable for the parameter $\kappa_{w}$ of any vertex $w$ (they are i.i.d.) and $K|\kappa$ is the random degree of a vertex of parameter $\kappa$ when all vertices are independently and uniformly distributed on the circle [10]. These choices for $R$ and $\mu$ are without lost of generality.
@@
$\mathbb{P}[\mathcal{G}=G|\theta,\kappa,\beta]$ $=\prod_{(u,v)\in E}\mathbb{P}[a_{uv}=1|\theta,\kappa,\beta]$
×$\hskip 10.0pt\times\prod_{(u,v)\not\in E}(1-\mathbb{P}[a_{uv}=1|\theta,\kappa,\beta])$
@@
where $\binom{V}{2}$ is the set of all combinations of two vertices, and $\beta_{uv}^{\pm}$ equals $\beta$ if $a_{uv}=1$ and equals $-\beta$ otherwise.
@@
The likelihood $\mathbb{P}[\mathcal{G}=G|\theta,\kappa,\beta]$ is given by Eq. (II.1.1). To complete the specification of Eq. (3), we use independent prior densities
@@
with $\beta_{0}=3$ and $\sigma=2$, and where $\mathds{1}[\cdot]$ is the indicator function (equals $1$ is its argument is true, and 00 otherwise). Since known networks tend to have a heavy-tailed degree distribution, we use a half-Cauchy prior
@@
except for the two vertices with the highest degree, whose position is restricted (the reason is technical and has to do with avoiding degeneracy in the posterior; see below). With these choices, the posterior distribution is proper and can express correlations between the parameters even if the priors are independent. The normalization constant $\mathbb{P}[\mathcal{G}=G]$ can be obtained by integration in principle, though it is not needed here; see the Sampling section below.
@@
Figure 2: Log-likelihood of the model when a single vertex is moved along the circle. Each dip is a divergence that occurs when a vertex is positioned at the same position as a disconnected vertex. These divergences form barriers in the objective function landscape (log-likelihood, posterior distribution), which is one of the reasons why hyperbolic embedding is a difficult problem. The likelihood is computed on the ground truth embedding of the synthetic graph of Fig. 1. Colors indicate the degree of the moved vertex (lowest degree in light blue, median degree in blue and highest degree in dark blue).
II.1.2 Symmetries and automorphisms
The $\mathbb{S}^{1}$ model exhibits fundamental symmetries that give rise to identifiability problems. While these symmetries are often overlooked when using maximization algorithms (which produce single solutions), they must be addressed in the Bayesian context. Indeed, our goal is to compute a distribution over embeddings, requiring us to distinguish between embeddings that differ merely by trivial reparametrizations (such as coordinate rotations) and those that represent genuinely distinct configurations. In this section, we examine the two classes of symmetries inherent in the model and present our approach for handling them.
@@ II.1.2 Symmetries and automorphisms
First, the model inherits the symmetries of the circle $\mathbb{S}^{1}$. Since the distance $d$ is invariant to rotations and reflections $\theta_{w}\mapsto-\theta_{w}\ \forall w$, the edge probabilities are preserved in the likelihood, and an infinite number of equivalent embeddings exist. Maximization algorithms break this symmetry at random, but we need to proceed with more caution when we move to a distribution over embedding. Without loss of generality, we fix the angular coordinate of the highest-degree vertex $u$ to $\theta_{u}=0$ to handle rotational symmetries. We remove the reflection symmetry by restricting the angular coordinate of the second highest-degree vertex $v$ to $\theta_{v}\in[0,\pi]$.
@@
Second, a graph is inherently labeled by its vertices, but any vertex relabelling that preserves the edges—an automorphism—leaves the likelihood invariant since the distances between connected and unconnected vertices are unchanged. For example, if a transformation $u\leftrightarrow v$ exists such that each $(u,w)\in E\Longleftrightarrow(v,w)\in E$ for any vertex $w$ other than $u$ and $v$, then we can also exchange $\kappa_{u}\leftrightarrow\kappa_{v}$ and $\theta_{u}\leftrightarrow\theta_{v}$ to preserve the likelihood.
The automorphisms cannot easily be avoided in the model and while the restriction on the highest-degree vertices helps with sampling, it does not guarantee an optimal compatibility between the embeddings. For instance, fixing $\theta_{u}=0$ implies no uncertainty on $\theta_{u}$ and restricting some $v$ to $\theta_{v}\in[0,\pi]$ can affect $|V|-2$ vertices instead of only $\theta_{v}$. To have the most coherent embedding sample, we opt for a post-processing alignment where we minimize the sum of squared angular separations to some arbitrary reference embedding over the automorphisms found using Nauty [72], the two possible reflections (reflection or no reflection) and the rotations $[0,2\pi)$.
In what follows, we will compare the embeddings obtained using this Bayesian framework with the ones obtained using Mercator [36]—a coordinate ascent algorithm that maximizes the Eq. (II.1.1). When embedding synthetic graphs generated with the $\mathbb{S}^{1}$ model, we will also use the original input parameters $(\theta,\kappa,\beta)$ as the “ground truth” embedding.
II.2 Sampling
With the model in place, we can now generate embeddings by calculating expectations over the posterior density of Eq. (3). For example, the expected position of a vertex can be calculated as
[…]
where $\Theta_{u}$ is the random variable of the angular position of vertex $u$, and similar integrals can be defined for all quantities of interests, like confidence interval for the positions. However, these integrals have, as far as we can tell, no closed-form solution, and we thus turn to sampling approximations computed using Markov Chain Monte Carlo (MCMC) algorithms. These algorithms tend to be slow (see Supplementary Note 1 for a discussion), but we show in the Multimodality section below that there is evidence for their necessity to adequately tackle this problem.
@@ II.2 Sampling
Unfortunately, we have found that standard MCMC algorithms are not well-adapted to the $\mathbb{S}^{1}$ model. Indeed, for most graphs $G$, the posterior distribution has a complicated geometry defined by several steep barriers that separate the posterior distribution in a multitude of small regions. This is the case even for a small graph of $30$ vertices, as Fig. 2 shows: Changing the embedding coordinate of a single vertex reveals multiple local maxima of the likelihood (and thus posterior density). The model predicts that two vertices with nearby embedding coordinates should be connected with high probability. In the extreme case where two disconnected vertices have the same angular coordinate, the likelihood approaches 00, causing the sharp dips shown in the Figure. Furthermore, the gradients are undefined due to their dependence on the distance function in Eq. (1). Since the posterior geometry is filled with these barriers for sparse and large graphs (the number of modes increases with the number of disconnected pairs of vertices), hyperbolic embedding can be challenging for standard MCMC algorithms.
In what follows, we first shows that both a simple random walk algorithm and a sophisticated gradient-based algorithm cannot adequately sample the posterior of a small graph as a result. Then, we demonstrate that our proposed algorithm, BIGUE, can avoid this problem by supplementing cluster-based transformations to a random walk.
@@
Figure 3: Statistics of Markov chain Monte Carlo algorithms when embedding the synthetic graph of Fig. 1. Results for the random walk algorithm (RW) and for our method (BIGUE) that uses both cluster transformations and a random walk (RW+CT) are displayed in red and blue, respectively. The orange curves show the results for dynamic Hamiltonian Monte Carlo (HMC) for the differentiable $\mathbb{S}^{1}$ model described in Supplementary Note 3. The green line displays the maximum log-likelihood obtained from Mercator after $10$ runs, each with $10$ refinements [36] and the black line is the ground truth embedding’s log-likelihood. (a) Normalized autocovariance averaged over all parameters and chains at different lags. (b-c) Traceplot of log-likelihood of a simulated Markov chain initialized without (panel b) and with (panel c) access to ground truth information. When the ground truth is unknown, we initialize $\kappa_{u}=\deg(u)$ and draw the other parameters from their priors. (d) Traceplot of log-likelihood of a Markov chain after thinning the chain shown in panel b. The blue line of this panel is the sample shown in Fig. 1. In each case, we compute $\hat{R}_{\text{max}}$, the maximum potential scale reduction factor for all parameters after the iterations removed from the warm-up (grey part of traceplots)—values close to 1 are desirable. Four independent chains were simulated to compute $\hat{R}$ (shown on each panel with the color corresponding to the sampling algorithm) and the autocovariance, but only one is displayed (representative of the four).
II.2.1 Baseline algorithms